# Revisiting Koehler’s experiment of measuring the ratio of the specific heats of air by self-sustained oscillations: a more concise theoretical interpretation

### v1

@@ IV.1 Difficulties in analyzing periodic solutions of the 3D PWLD system
Let us first provide a detailed explanation of the difficulty in rigorously analyzing the periodic solutions of system (10). The Poincaré mapping is a powerful tool for studying periodic solutions in dynamical systems. However, for general PWLD systems, constructing a Poincaré map requires careful treatment of transitions across the switching surface within each subspace, which presents significant challenges. For system (10), we treat the separation plane $\{y=0\}$ as the Poincaré section. As discussed previously, when the system operates in any half-space, the velocity of the ball can be represented as
@@
where $\alpha=\frac{G}{2}$, $\Omega=\sqrt{1-\frac{G^{2}}{4}}$. According to Eq. (9a), the expression for the coordinates of the ball can be obtained as
@@
By solving $y(\tau)=0$, in principle, the state of the system when switching between subspaces and its running time in each subspace can be obtained. However, $y(\tau)=0$ is too complex to yield analytical solutions, which complicates the construction of Poincaré mappings to analyze the periodic solution behavior of the system.
IV.2 Simplified problem: 2D PWLD system with one separation straight line
From the analysis above, it is evident that the trajectory of the system in any half-space, when projected onto the $(u,p)$ plane, appears as a simple spiral around the focus. However, the switching between subspaces is controlled by the condition $y=0$ in the third dimension, which presents significant challenges in solving for the switching state. To avoid this difficulty, we simplify system (10) by reducing its dimensionality, resulting in a purely 2D PWLD system as follows:
$\dot{\mathbf{x}}=\begin{cases}M_{+}\mathbf{x}+\mathbf{e}_{+}&\mathrm{if}\quad p\leq 0\\ M_{-}\mathbf{x}+\mathbf{e}_{-}&\mathrm{if}\quad p>0\end{cases}$ (16)
@@ IV.2 Simplified problem: 2D PWLD system with one separation straight line
$\mathbf{x}=\begin{pmatrix}u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{-}\end{pmatrix}$,
@@
$u^{*}_{+}=A\left[J-G_{+}(p_{0}-1)\right]<0$ (the ball’s motion tends towards the steady state of falling),
$u^{*}_{-}=A\left[J-G_{-}(p_{0}-1)\right]>0$ (the ball’s motion tends towards the steady state of rising).
Figure 2: (a) and (b) The phase portrait and transition map of 2D system with separation line $\{p=0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The red solid circle and blue hollow circle are two foci, while the hollow upward triangles and downward triangles are the intersections of the trajectory and the separation line. (c)-(f) The oscillations of $u$ and $p$, respectively. (d), (f) correspond to the enlarged display of shadow region in (c), (e) respectively.
@@
The periodic solutions of system (16) can be intuitively understood. Figure 2 illustrates a typical periodic solution, with parameters $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. For all simulations presented in this paper, the values of $G_{\pm}$ remain the same. Fig. 2(a) and 2(b) show its phase portrait and the corresponding transition map of the periodic steady state, respectively. In each half-plane, the trajectory is well-defined, spiraling clockwise around a focus. At a given moment, the trajectory intersects the separation line $p=0$ at the point $(u_{1},0)$, as shown in Fig. 2(b). The trajectory then enters the half-plane $\{p\leq 0\}$, continuing its clockwise spiral while asymptotically approaching the focus $(u_{+}^{*},0)$. It intersects the line $p=0$ again at $(u_{2},0)$. During this phase, the rotation radius relative to focus $(u_{+}^{*},0)$, which corresponds to the amplitude of the oscillation, decays exponentially, from a length of $(u_{1}-u_{+}^{*})$ to $(u_{+}^{*}-u_{2})$.
@@
Subsequently, the system enters the half-plane $\{p>0\}$, switching its focus to the point $(u_{-}^{*},0)$ and continuing its spiral towards this new focus. Since $(u_{-}^{*},0)$ is located to the right of $(u_{+}^{*},0)$, the radial distance $(u_{-}^{*}-u_{2})$ is greater than $(u_{+}^{*}-u_{2})$. Thus, when the system switches subspaces, the rotation radius increases abruptly, compensating for the previous decay. Starting with $(u_{-}^{*}-u_{2})$ as the new rotation radius, the trajectory spirals clockwise around $(u_{-}^{*},0)$ until it intersects the line $p=0$ once more. For the system to maintain a periodic steady state, the trajectory must return to the point $(u_{1},0)$, completing one full cycle of motion.
@@
Remarkably, $p=0$ is the line connecting the two foci. From a geometric perspective, it is evident that the trajectory in either half-plane rotates through an angle of $\pi$, as shown in Fig. 2(b). Specifically, the system spends $\frac{T_{\pm}}{2}$ in each half-plane, where $T_{\pm}=\frac{2\pi}{\Omega_{\pm}}\sim 2\pi$, with $\Omega_{\pm}=\sqrt{1-\frac{G_{\pm}^{2}}{4}}\sim 1$ for $G_{\pm}\ll 1$. Therefore, the overall period of the system’s complete motion remains approximately $2\pi$, which is near the reduced Rüchardt period $T_{0}=2\pi$. Having established the running time in each subspace, we now proceed to calculate the amplitude corresponding to the periodic solution as follows:
@@
$u_{1}=u_{-}^{*}+(u_{-}^{*}-u_{2})e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}},$
$u_{1}=\frac{u_{-}^{*}(1+e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}})-u_{+}^{*}e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}}(1+e^{-\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}})}{1-e^{-(\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}+\frac{G_{-}}{2}\cdot\frac{T_{-}}{2})}}\sim\frac{2(u_{-}^{*}-u_{+}^{*})}{\pi(G_{+}+G_{-})},\quad if\quad G_{\pm}\ll 1$ (18)
@@
$\frac{T_{+}}{2}=3.1775,\quad\frac{T_{-}}{2}=3.1438,\quad T=6.3214,$ (19)
@@
$u_{1}=11.1352,\quad u_{2}=-10.7759.$ (20)
These results agree exactly with the numerical simulation shown in Fig. 2(d).
According to Eq.(18) and Eqs.(17), when $G_{\pm}\ll 1$, the rotation radius of the trajectory can be approximated as $\sim\frac{u_{1}-u_{2}}{2}\sim u_{1}-u_{+}^{*}>u_{1}\gg u_{-}^{*}-u_{+}^{*}$. Even if the separation line (i.e., $p=p_{1}\neq 0$}) is not the line determined by the two foci, the conclusion that the trajectory’s average rotation radius is significantly larger than the distance between the two foci remains valid. This indicates that, over one period, the trajectory can be approximated as a complete rotation around an equivalent point, corresponding to an angle of $2\pi$. Consequently, the system’s period remains close to the Rüchardt period, although the deviation becomes slightly more pronounced as $p_{1}\neq 0$.
For instance, Figure 3 illustrates the periodic solution when the separation line is given by $p=0.003$. Within a complete cycle, the system resides in the half-planes $p\leq 0$} and p>0{p>0} for durations of $3.9815$ and $2.4104$, respectively, with corresponding rotation angles of $1.2530~\pi$ and $0.7667~\pi$. The total rotation angle for the cycle is approximately $2.0197~\pi$, demonstrating that the trajectory still maintains a near-complete $2\pi$ rotation despite the shifted separation line.
Figure 3: (a) and (b) The phase portrait and transition map of 2D system with separation line $\{p=p_{1}\neq 0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The red solid circle and blue hollow circle are two foci, while the hollow upward triangles and downward triangles are the intersections of the trajectory and the separation line. (c) The oscillations of $u$. (d) The enlarged display of shadow region in (c).
@@ IV.3 The original 3D PWLD system with one separation plane y=0
Figure 4: (a) and (b) The phase portrait and transition map of 3D system with separation plane $\{y=0\}$. The parameters of numerical simulation are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. Under the setting of these parameters, the oscillation of displacement $y$ is symmetrical with respect to plane $\{y=0\}$. (c)-(h) The oscillations of $y$, $u$ and $p$, respectively. (d), (f) and (h) correspond to the enlarged display of shadow region in (c), (e) and (g) respectively.
@@
Figure 5: The phase portrait, transition map and $y$, $u$ oscillation of 3D system with separation plane $\{y=0\}$. The simulation parameters corresponding to the figures in the left column are $G_{+}=0.3$, $G_{-}=0.075$, and $J=1.2\times 10^{-3}$. The simulation parameters corresponding to the figures in the right column are $G_{+}=0.3$, $G_{-}=0.075$, and $J=0.9\times 10^{-3}$. under the setting of these parameters, the oscillation of displacement $y$ is asymmetrical with respect to plane $\{y=0\}$.
@@
Figures 4 and 5 present the numerical simulation results of the 3D system (10) under different $J$ (gas pumping). Figures 4 shows the case of symmetric oscillation of the displacement $y$ of the small ball relative to the position of the hole. When the ball passes through the small hole position, the pressure inside the container approaches the equilibrium pressure $P_{0}$, corresponding to the dimensionless reduced pressure $p\sim 0$. As seen in Fig. 4(b), in the projection plane, the equivalent separation line nearly coincides with the line connecting the two foci, corresponding to the motion depicted in Fig. 2. Over a complete period, the trajectory rotates through angles of $0.9921~\pi$ and $0.9996~\pi$ in the two half-planes, respectively. The corresponding running time 3.1524 and 3.1424 are very close to $\frac{T_{+}}{2}~(3.1778)$ and $\frac{T_{-}}{2}~(3.1438)$, respectively, and the entire period $T=6.2948$ is very close to the reduced Rüchardt period of $2\pi$. Additionally, the values of $u_{1}$ (11.1352) and $u_{2}$ (-10.7759) calculated by Eq.(20) are also very close to those in Fig. 4 (f) (11.1144 and -10.7532).
@@
Figures 5 illustrates two cases of asymmetric oscillation of the displacement $y$. In these cases, the equivalent separation line deviates from the line connecting the two foci, corresponding to the motion depicted in Fig. 3. Over a complete period, the trajectory no longer rotates through angles close to $\pi$ in each half-plane. However, the total angular displacement still approaches $2\pi$, and the entire period (6.305 or 6.2841) remains close to the reduced Rüchardt period of $2\pi$.
@@
In Koehler’s experiment, it is essential to adjust the gas flow rate to ensure that the small ball oscillates symmetrically about the position of the hole. In the theoretical analysis presented in Koehler’s paper Koehler1950 , it is assumed that the small ball spends equal amounts of time above and below the hole during each oscillation cycle. This assumption is necessary for carrying out the integration required for Fourier series analysis. The analysis in our paper offers a clearer understanding of the rationale behind this requirement, avoiding the need for complex mathematical calculations.
@@ V Conclusion
In this paper, we revisit Koehler’s experiment and provide a concise theoretical interpretation of the self-sustained oscillations observed in the system. Without resorting to complex calculations, our analysis demonstrates that the oscillation frequency in Koehler’s experiment approximates the Rüchardt frequency. After explicitly presenting the governing equations of the model, we found that it is challenging to mathematically determine the precise parameter conditions under which the system exhibits periodic solutions close to the Rüchardt frequency. However, we prove that if such periodic solutions exist, specifically, solutions in which the small ball passes through the hole position once upward and once downward during a single period, then the system’s motion frequency is indeed very close to the Rüchardt frequency.
The self-sustained oscillations in Koehler’s experiment differ fundamentally from simple harmonic or linearly forced oscillations, representing an early nonlinear phenomenon typically encountered by undergraduates. While students can complete the experimental procedures without distinguishing between the principles underlying Koehler’s and the Rüchardt experiments, we believe that for students proficient in observation and inquiry, exploring the theoretical issues behind Koehler’s experiment can deepen their understanding of periodic motion and offer a practical scenario for studying dynamical systems.

### v2

@@ IV.1 Difficulties in analyzing periodic solutions of the 3D PWLD system
We first examine the difficulty of rigorously analyzing the periodic solutions of system (11). Poincaré mappingstrogatz2024nonlinear is a powerful tool for studying periodic solutions in dynamical systems. However, for general PWLD systems, constructing a Poincaré map requires careful treatment of transitions across the switching surface within each subspace, which presents significant challenges. For system (11), we treat the separation plane $\{y=0\}$ as the Poincaré section, which is a hypersurface transversally crossed by the system trajectories such that successive intersections define a discrete return map. In the present system, trajectories cross this plane with nonzero velocity in the $y$ direction, ensuring that the induced Poincaré map is well-defined. As discussed previously, when the system operates in either half-space, the velocity of the ball can be represented as
@@
where $\alpha=G/2$, $\Omega=\sqrt{1-G^{2}/4}$. According to Eq. (10a), the coordinates of the ball can be obtained as
@@
By solving $y(\tau)=0$, in principle, the state of the system when switching between subspaces and its running time in each subspace can be obtained. However, the condition $y(\tau)=0$ is generally too complex to yield closed-form analytical solutions, which complicates the construction of Poincaré maps for analyzing periodic solutions. Although such maps can be constructed numerically, this typically requires case-by-case computation, limiting the possibility of obtaining general analytical insight.
IV.2 Simplified problem: 2D PWLD system with one separation line
From the analysis above, it is evident that the system’s trajectory in either half-space, when projected onto the $(u,p)$ plane, appears as a simple spiral around the focus. To avoid the issue of switching between subspaces, we reduce system (11) to a 2D PWLD system:
$\dot{\mathbf{z}}=\begin{cases}M_{+}\mathbf{z}+\mathbf{e}_{+}&\mathrm{if}\quad p\leq 0\\ M_{-}\mathbf{z}+\mathbf{e}_{-}&\mathrm{if}\quad p>0\end{cases}$ (18)
@@ IV.2 Simplified problem: 2D PWLD system with one separation line
$\mathbf{z}=\begin{pmatrix}u\\ p\end{pmatrix},\quad M_{+}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{+}\end{pmatrix},\quad M_{-}=\begin{pmatrix}0&A\\ -\frac{1}{A}&-G_{-}\end{pmatrix}$,
@@
$u^{*}_{+}=A\left[J-G_{+}(p_{0}-1)\right]<0,\qquad u^{*}_{-}=A\left[J-G_{-}(p_{0}-1)\right]>0.$
Figure 2: (Color online) (a) Phase portrait of the 2D system with separation line $\{p=0\}$, showing the full dynamics including the transient regime (left), and the resulting periodic steady-state limit cycle with trajectory switching across the separation line (right). The numerical parameters are $A=1782$, $G_{+}=0.225$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The red solid and blue hollow circles denote the two fixed points, while the hollow upward and downward triangles indicate the intersections of the trajectory with the separation line. (b)-(c) Oscillations of $u$ and $p$, respectively. The right-hand portions show enlarged views of the shaded regions.
@@
The periodic solutions of system (18) can be intuitively understood. Figure 2 illustrates a typical periodic solution, with parameters $A=1782$, $G_{+}=0.225$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. In all numerical solutions presented in this paper, the values of $G_{\pm}$ and $A$ are kept fixed; these values are based on typical experimental estimates (see Appendix B). Figure 2(a) shows the phase portrait of the full dynamics including the transient regime (left), and the resulting periodic steady-state limit cycle, illustrating the trajectory switching between subspaces (right). In each half-plane, the trajectory is well-defined, spiraling clockwise around a focus. When the trajectory in $p>0$ intersects the separation line $p=0$ at point $(u_{1},0)$, it then enters the half-plane $\{p\leq 0\}$, continuing its clockwise spiral while asymptotically approaching the focus $(u_{+}^{*},0)$. It intersects the line $p=0$ again at $(u_{2},0)$. During this phase, the rotation radius, defined as the instantaneous Euclidean distance in phase space between the trajectory point $(u(t),p(t))$ and the fixed point $(u_{+}^{*},0)$, decays exponentially from $(u_{1}-u_{+}^{*})$ to $(u_{+}^{*}-u_{2})$. This quantity corresponds to the oscillation amplitude.
@@
Subsequently, the system enters the half-plane $\{p>0\}$, switching its focus to the point $(u_{-}^{*},0)$ and continuing its spiral towards this new focus. Since $u_{-}^{*}<u_{+}^{*}$, the distance $(u_{-}^{*}-u_{2})$ is greater than $(u_{+}^{*}-u_{2})$. Thus, when the system switches subspaces, the rotation radius increases abruptly, compensating for the previous decay. Starting with $(u_{-}^{*}-u_{2})$ as the new rotation radius, the trajectory spirals clockwise around $(u_{-}^{*},0)$ until it intersects the line $p=0$ once more. For the system to maintain a periodic steady state, the trajectory must return to the point $(u_{1},0)$, completing one full cycle of motion.
@@
Remarkably, $p=0$ is the line connecting the two foci. From a geometric perspective, it is evident that the trajectory in either half-plane rotates through an angle of $\pi$, as shown in Fig. 2(a). Specifically, the system spends $T_{\pm}/2$ in each half-plane, where $T_{\pm}=2\pi/\Omega_{\pm}\sim 2\pi$, with $\Omega_{\pm}=\sqrt{1-G_{\pm}^{2}/4}\sim 1$ for $G_{\pm}\ll 1$. Therefore, the nondimensional period of the motion remains approximately $2\pi$, equal to the reduced Rüchardt period $T_{0}$. Having established the running time in each subspace, we now proceed to calculate the amplitude corresponding to the periodic solution as follows:
@@
$u_{1}=u_{-}^{*}+(u_{-}^{*}-u_{2})e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}}.$
In the limit that $G_{\pm}\ll 1$,
$u_{1}=\frac{u_{-}^{*}(1+e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}})-u_{+}^{*}e^{-\frac{G_{-}}{2}\cdot\frac{T_{-}}{2}}(1+e^{-\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}})}{1-e^{-(\frac{G_{+}}{2}\cdot\frac{T_{+}}{2}+\frac{G_{-}}{2}\cdot\frac{T_{-}}{2})}}\sim\frac{2(u_{-}^{*}-u_{+}^{*})}{\pi(G_{+}+G_{-})}.$ (20)
@@
$\frac{T_{+}}{2}=3.18,\quad\frac{T_{-}}{2}=3.14,\quad T=6.32,$ (21)
@@
$u_{1}=11.14,\quad u_{2}=-10.78.$ (22)
These results agree exactly with the numerical simulation shown in Fig. 2(b).
According to Eq.(20) and Eqs.(19), when $G_{\pm}\ll 1$, the rotation radius of the trajectory can be approximated as $(u_{1}-u_{2})/2\sim u_{1}-u_{+}^{*}>u_{1}\gg u_{-}^{*}-u_{+}^{*}$. Even if the separation line is not the one passing through the two foci (i.e., $p=p_{1}\neq 0$}), the conclusion that the trajectory’s average rotation radius is significantly larger than the distance between the two foci remains valid. This indicates that, over one period, the trajectory can be approximated by a complete rotation around an effective center located between the two foci, corresponding to an angle of $2\pi$. Consequently, the system’s period remains close to the Rüchardt period, although the deviation becomes slightly more pronounced as $p_{1}$ moves away from zero.
For instance, Figure 3 illustrates the periodic solution when the separation line is given by $p=0.002$. Within a complete cycle, the system resides in the half-planes $p\leq 0$} and p>0{p>0} for durations of $3.98$ and $2.41$, respectively, with corresponding rotation angles of $1.25~\pi$ and $0.77\pi$. The total rotation angle for the cycle is approximately $2.02~\pi$, demonstrating that the trajectory still maintains a near-complete $2\pi$ rotation despite the shifted separation line.
Figure 3: (Color online) (a) Phase portrait of the 2D system with separation line $\{p=0.002\}$, showing the full dynamics including the transient regime (left), and the resulting periodic steady-state limit cycle with trajectory switching across the separation line (right). All other parameters and graphical conventions are the same as in Fig. 2. (b) Oscillations of $u$. The right-hand portion shows an enlarged view of the shaded region.
@@ IV.3 The original 3D PWLD system with one separation plane y=0
Figure 4: (Color online) (a) Projected phase portrait of the 3D system with separation plane $\{y=0\}$. The numerical parameters are $A=1782$, $G_{+}=0.225$, $G_{-}=0.075$, and $J=1.0628\times 10^{-3}$. The hollow upward and downward triangles denote the projections onto the $(u,p)$ plane of the intersections where the trajectory crosses the separation plane $\{y=0\}$ upward and downward, respectively. As the trajectory approaches the steady-state periodic orbit, these two sequences of points converge to $u_{1}$ and $u_{2}$. (b) Oscillations of the displacement $y$, showing symmetry about the plane $\{y=0\}$. The right-hand portion shows an enlarged view of the shaded region.
@@
Figure 5: (Color online) Each column contains, from top to bottom, the projected phase portrait of the 3D system with separation plane $\{y=0\}$ and the oscillations of the displacement $y$. The parameters are $A=1782$, $G_{+}=0.225$ and $G_{-}=0.075$. For the left column, $J=1.2\times 10^{-3}$, while for the right column $J=0.9\times 10^{-3}$. For these parameter sets, the displacement $y$ exhibits asymmetric oscillations with respect to the plane $\{y=0\}$.
@@
Figures 4 and 5 present the numerical results of the 3D system (11) under different $J$ (pumping). Figure 4 shows the case where the oscillation is symmetric about the plane $\{y=0\}$. When the ball passes through the small hole position, the pressure inside the container approaches the equilibrium pressure $P_{0}$, corresponding to the dimensionless reduced pressure $p\sim 0$. As shown in Fig. 4(a), the equivalent separation line in the projected $(u,p)$ plane nearly coincides with the line connecting the two foci, resembling the motion shown in Fig. 2. Over a complete period, the trajectory rotates through angles of $0.9921~\pi$ and $0.9996~\pi$ in the two half-planes, respectively. The corresponding running time 3.15 and 3.14 are very close to $\frac{T_{+}}{2}~(3.18)$ and $\frac{T_{-}}{2}~(3.14)$, respectively, and the entire period $T=6.29$ is very close to the reduced Rüchardt period of $2\pi$.
@@
Figure 5 illustrates two cases of asymmetric oscillation of the displacement $y$ about the plane $\{y=0\}$. In these cases, the equivalent separation line deviates from the line connecting the two foci, resembling the motion shown in Fig. 3. Over a complete period, the trajectory no longer rotates through angles close to $\pi$ in each half-plane. However, the total angular displacement still approaches $2\pi$.
@@
In Koehler’s experiment, it is essential to adjust the gas flow rate to ensure that the small ball oscillates symmetrically about the position of the hole. In the theoretical analysis presented in Koehler’s paper Koehler1950 , it is assumed that the small ball spends equal amounts of time above and below the hole during each oscillation cycle. This assumption is necessary for carrying out the integration required for Fourier series analysis. The analysis in our paper clarifies the physical meaning of this requirement, showing that the symmetry about $y=0$ is not a strict requirement, but rather a limiting condition: the closer the motion is to being symmetric about $y=0$, the closer the oscillation period approaches the ideal Rüchardt period.
@@ V Conclusion
Koehler’s experiment provides a simple yet instructive realization of self-sustained nonlinear oscillations in a gas–mechanical system. Our analysis shows that, under conditions where the motion remains periodic and the small ball crosses the hole position once in each direction during a cycle, the resulting oscillation frequency is close to the classical Rüchardt frequency.
The measurement of $\gamma$ is typically introduced in lower-level undergraduate laboratory courses. The nonlinear dynamical concepts involved in the present work may therefore present a certain mathematical barrier for many students. For this reason, in practical teaching, even when discussing Koehler’s experiment, it is useful to first adopt the physical framework underlying the Rüchardt experiment as the primary pedagogical model. After establishing this baseline understanding, the differences between the two experimental configurations can be highlighted. More advanced students can be encouraged to reproduce and extend the analysis presented in this work. This approach may help deepen their understanding of periodic motion and provide a concrete context for studying dynamical systems.
Finally, when multiple self-sustained oscillators are coupled, synchronization phenomena may emerge, as exemplified by the well-known synchronization of metronomes10.1119/1.1501118 . Introducing coupling between multiple Koehler apparatuses to demonstrate synchronization would be an interesting extension. In contrast to metronomes, whose operation is limited by a finite energy reservoir (the spring-driven mechanism), the oscillations in the Koehler system can, in principle, be sustained for an arbitrarily long time under continuous driving, making it a potentially robust platform for studying synchronization phenomena.
Acknowledgments
This work was funded by the Research Project Supported by Shanxi Scholarship Council of China (Grant No. 2023-033) and by the Fundamental Research Program of Shanxi Province (Grant No. 202303021221071).